# A simple fourth order propagator based on the Magnus expansion in the Liouville space: Application to a $\Lambda$-system and assessment of the rotating wave approximation

### v3

@@ III.4 Comparison of different fourth-order methods
It is important to assess the accuracy and stability of our simple 4th order propagators compared to other methods often used in quantum dynamics. To this end, we compared ours with two widely used 4th order methods,[63, 64] explicit 4th order Runge-Kutta (RK4) method[63] and implicit 4th order Adams-Moulton (AM4) method,[64] We also implemented an optimized 4th order CFME (O4-CFME)[41, 36] that utilizes three exponentials and Gauss-Legendre points. All of these are summarized in Appendix F. The error of each method was determined according to
@@
where our 6th order ME-based propagator[29] was used to calculate the reference propagator $\mathcal{U}_{ref}$. Errors for both a short time ($t=2\pi$) and a long time ($t=16\pi$) were calculated for each case.
The results for the closed-system dynamics B-I are plotted in Fig. S10. We see that our 4th order ME and CFME propagators have about an order of better accuracy than the RK4 and AM4 methods, both in the short-time and the long-time limits. However, our results are significantly less accurate than those of the O4-CFME method.[41, 36] This is understandable from the fact that the latter utilizes effectively one more time for each time step and one more exponential. In addition, the superior performance of O4-CFME suggests the benefit of using Gauss-Legendre points and at least three exponentials if possible.
The results for the open-system dynamics of B-II are shown in Fig. S11. While the short-time dynamics show that our ME-based methods work better than the RK4 and AK4 methods, its benefit worsens for the long-time dynamics. As a result, the accuracy of our 4th order ME propagator with commutator becomes comparable to the RK4 method, whereas our CFME method remains more accurate than the latter. On the other hand, the performance of the O4-CFME method[41, 36] remains best even in this case. Similar behavior can be seen for all models of open-system dynamics considered in this work with some exceptions (see the SM).
@@
Figure 10: Comparison of 4th order ME-based methods with the RK4 and AM4 methods for the closed-system dynamics of case B-I (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19).
Figure 11: Comparison of 4th order ME-based methods with the RK4 and AM4 methods for the open-system dynamics of case B-II (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19)
@@ IV Conclusion
The main result of our work presented here is the extension and test of our simple 4th order ME-based propagator to the Liouville space for both closed and open system dynamics, as represented by Eqs. (3), (8), (11), and (12). These propagators are simple to use and can be naturally incorporated into broad range of quantum dynamics calculations while based on evaluation of Hamiltonians at a mid point in addition to end points of each time step. In particular, our 4th order CFME, Eq. (12) was confirmed to be consistently better than the RK4 and AM4 methods while being simple to implement.
@@
Another important aspect of our work is detailed assessment of the RWA, which has been used widely due to its convenience and conceptual clarity, and is indeed well established as an accurate approximation for conventional spectroscopy measurements in the weak field limit. However, beyond that limit and/or in the presence of multiple pulses, its reliability remains poorly understood. Considering recent experimental advances in nonlinear spectroscopy,[65] QC, and QS and the need for more accurate calculation methods of optical signals involved, better understanding of how the RWA breaks down is an important issue. Results of our calculation for a prototypical $\Lambda$-system are consistent with the general notion that the RWA becomes reliable in the weak field and steady state limits. However, for moderate control and probe field strengths, we find some of the results based on the RWA Hamiltonian exhibit subtle deviations from those of full Hamiltonian. For a strong control field strength, there are substantial differences even at the qualitative level even though the probe field remains moderate. Although these have been expected, the actual quantitative comparison can help better assessment of the RWA in general.
@@
Our numerical tests including the O4-CFME method[41, 36] demonstrate superior performance of the ME-based approaches and the benefit of using more Gauss-Legendre points and more exponentials instead of commutators. As yet, these results are limited to the $\Lambda$-model with periodic time dependences as considered here. Thus, further tests are needed for other systems and more complex time dependences in order to establish a reliable general trends of performance.
@@
It is also noteworthy to mention the stability of our ME based numerical method for the non-Hermitian dynamics (e.g. Figs. 4). We found that the trace of unity is maintained with only negligible deviations (on the order of $10^{-12}$ in the worst case) due to numerical error, even in the long-time limit. This suggests that, for the time-dependent system Hamiltonian $\hat{H}(t)$ with time-independent Lindblad dissipative terms as in Eq. (15), the ME truncated at any order produces a trace preserving map. In fact, we find numerically that it produces CPTP maps, which not only guarantee that the trace is maintained but ensure that the resulting density operator is physical even if it only describes a subsystem. An analytical proof for the closed system case is shown in Appendix B. For the open system case, the CPTP map is produced for small enough time step sizes, which ensures that the first order term of the ME is larger than the subsequent terms. This can be understood from the fact that the first order term of the ME defines a CPTP map, while it has been shown that higher order terms contribute to breaking the complete positivity.[37]
@@ Acknowledgements.
S.J.J. acknowledges major support of this research from the US Department of Energy, Office of Sciences, Office of Basic Energy Sciences (DE-SC0021413; DE-SC0026114) and partial support during the initial stage of this research from the National Science Foundation (CHE-1900170). S.J.J. also acknowledges support from Korea Institute for Advanced Study through its KIAS Scholar program for the visit during summer.
@@
Appendix A Liouville-space transformation
@@
Appendix B Proof that truncated Magnus expansion produces a completely positive trace preserving (CPTP) map
Using the same Liouville space notation as in Appendix A, a map is CPTP if and only if it can be written in the following form:[67]
@@ Appendix B Proof that truncated Magnus expansion produces a completely positive trace preserving (CPTP) map
where $\mathcal{M}_{n}$ is a functional of superoperators in the exponent of Eq. (6), consisting of iterated integrals and commutators (for precise definitions, see Ref. 29) . However, note that because $\hat{H}(t)\otimes\hat{I},\hat{I}\otimes\hat{H}(t)^{T}]=0$, it becomes possible to separate the exponent in Eq. (B) into two terms:
@@
Appendix C Closed system dynamics in the rotating wave frame (RWF)
@@
Appendix D Solution of closed system dynamics using RWA at the TPR conditions
@@
Appendix E Dynamics for periodic Hamiltonian
@@
Appendix F Other 4th order methods
F.1 Runge-Kutta and Adams-Moulton methods
We here describe the 4th order Runge-Kutta (RK4)[63] and Adams-Moulton (AM4)[64] numerical methods for solving a differential equation of the following generic form:
$\frac{dy(t)}{dt}=f(t,y(t))=A(t)y(t).$ (40)
For the case of dynamics in the Hilbert space, $f(t,y(t))=(-i\hat{H}(t)/\hbar)y(t)$, with $y(t)$ the time evolution operator. For the case of the dynamics in the Liouville space, $f(t,y(t))=-{\mathcal{L}}(t)y(t)$, with $y(t)$ the density operator. This notation is adopted for easier comparison with the literature on ordinary differential equations. [63, 64]
@@ F.1 Runge-Kutta and Adams-Moulton methods
F.2 Optimized commutator free Magnus Expansion
@@
In the above equation, $A_{i}$ is shorthand for $A(t+\delta tc_{i})$ where $c_{i}$ is the ith Gauss-Legendre point defined as $c_{1}=\frac{1-\sqrt{3/5}}{2}$ , $c_{2}=\frac{1}{2}$ , $c_{3}=\frac{1+\sqrt{3/5}}{2}$, and the coefficients $a_{ij}$ are defined as the matrix elements of $a$ below:
@@
G.1 Some numerical results for cases in Table I of main text
Results of calculation for Cases A-I, A-II, C-I, and C-II in Table I of the main text are provided here. All Python code to reproduce data and figures can be found at https://github.com/TanerTure/floquet_dynamics.
Figure S2: Elements of the time dependent system density operator for the case A-I in the Table I of the main text, a closed system unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (8) in the main text.
Figure S3: Elements of the time dependent system density operator for the case A-II in Table I of the main text, an open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (11) in the main text.
Figure S4: Elements of the time dependent system density operator for the case C-I in Table I of the main text, a closed system unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (8) in the main text.
Figure S5: Elements of the time dependent system density operator for case C-II in Table I of the main text, an open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (11) in the main text.
Figure S6: Elements of the time dependent system density operator in the long time limit for the case A-II in Table I of the main text, an open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (11) in the main text.
Figure S7: Elements of the time dependent system density operator in the long time limit for the case B-II in Table I of the main text, an open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (11) in the main text.
Figure S8: Elements of the time dependent system density operator in the long time limit for the case C-II in Table I of the main text, an open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (11) in the main text.
(a) (b)
(c) (d)
Figure S9: Steady state limits of populations at state 2 ($\rho_{22}$) for the open system of case A-II with full Hamiltonian (a) and RWA (b), for which times required for convergence are respectively plotted in (c) and (d).
Figure S10: Comparison of fourth-order Magnus Expansion, Runge-Kutta, and Adams-Moulton methods for the closed-system dynamics of case A-I (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19) in the main text.
Figure S11: Comparison of fourth-order Magnus Expansion, Runge-Kutta, and Adams-Moulton methods for the open-system dynamics of case A-II (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19) in the main text.
Figure S12: Comparison of fourth-order Magnus Expansion, Runge-Kutta, and Adams-Moulton methods for the closed-system dynamics of case C-I (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19) in the main text.
Figure S13: Comparison of fourth-order Magnus Expansion, Runge-Kutta, and Adams-Moulton methods for the open-system dynamics of case C-II (a) for short time evolution ($t=2\pi$) and (b) for long time evolution $16\pi$. The error was calculated using Eq. (19) in the main text.
Figure S14: Plot of error (given by sixth-order - fourth-order) for the case B-I of the Table I, a unitary closed system dynamics, in the main text.
Figure S15: Plot of error (given by sixth-order - fourth-order) for the case B-II of the Table I, a nonunitary open system dynamics, in the main text.
Figure S16: Plot of error (given by sixth-order - fourth-order) for the case B-II of Table I, a nonunitary open system dynamics, in the steady state limit.

### v4

@@ III.4 Comparison of different fourth-order methods
In order to assess the accuracy and stability of our simple 4th order propagators, we compared ours with two widely used 4th order methods,[75, 76] explicit 4th order Runge-Kutta (RK4) method[75] and implicit 4th order Adams-Moulton (AM4) method,[76] We also implemented an optimized 4th order CFME (O4-CFME)[48, 43] that utilizes three exponentials and Gauss-Legendre points. Appendix G provides summaries of these three methods. The error of each method was quantified by the following definition of error:
@@
where $\mathcal{U}_{ref}$ represents calculation results based on our 6th order ME propagator.[36] Errors for both short time ($t=2\pi$) and long time ($t=16\pi$) dynamics were calculated for each algorithm.
Figure 10 shows results for the closed-system dynamics of case B-I. We see that our 4th order ME and CFME propagators have about an order of better accuracy than the RK4 and AM4 methods, both in the short-time and the long-time limits. However, our results are significantly less accurate than those of the O4-CFME method.[48, 43] This is understandable considering that the latter utilizes one more value of time for each step of integration and one more exponential. In addition, the superior performance of O4-CFME suggests the benefit of using Gauss-Legendre points and at least three exponentials if possible.
Figure 11 shows results for the open-system dynamics of case B-II. While the short-time dynamics show that our ME-based methods work better than the RK4 and AK4 methods, its benefit worsens for the long-time dynamics. As a result, the accuracy of our 4th order ME propagator with commutator becomes comparable to the RK4 method, whereas our CFME method remains more accurate than the latter. On the other hand, the performance of the O4-CFME method[48, 43] remains best even in this case. Similar behavior can be seen for all models of open-system dynamics considered in this work with some exceptions (see the SM).
@@
Figure 10: Comparison of 4th order ME-based methods with the RK4 and AM4 methods for the closed-system dynamics of case B-I (a) for short time evolution ($t=2\pi$) and (b) for long time evolution ($t=16\pi$). The error was calculated using Eq. (19).
Figure 11: Comparison of 4th order ME-based methods with the RK4 and AM4 methods for the open-system dynamics of case B-II (a) for short time evolution ($t=2\pi$) and (b) for long time evolution ($t=16\pi$). The error was calculated using Eq. (19)
@@ IV Conclusion
The main result of this work is the extension and test of our simple 4th order ME-based propagator to the Liouville space for both closed and open system dynamics, as represented by Eqs. (3), (8), (11), and (12). These propagators are simple to use and can be naturally incorporated into broad range of quantum dynamics calculations while based on evaluation of Hamiltonians at a mid point in addition to end points of each time step. In particular, our 4th order CFME, Eq. (12) was confirmed to be consistently better than the RK4 and AM4 methods while being simple to implement.
@@
Another important aspect of our work is detailed assessment of the RWA, which has been used widely due to its convenience and conceptual clarity, and is indeed well established as an accurate approximation for conventional spectroscopy measurements in the weak field limit. However, beyond that limit and/or in the presence of multiple pulses, its reliability remains poorly understood. Considering recent experimental advances in nonlinear spectroscopy,[77] QC, and QS and the need for more accurate calculation methods of optical signals involved, better understanding of how the RWA breaks down is an important issue. Results of our calculation for a prototypical $\Lambda$-system are consistent with the general notion that the RWA is reliable in the weak field and steady state limits. However, for moderate control and probe field strengths, we find some of the results based on the RWA Hamiltonian exhibit subtle deviations from those of full Hamiltonian. For a strong control field strength, there are substantial differences at the qualitative level even though the probe field remains moderate. Although these have been expected, the actual quantitative comparison provided here can help better assessment of the RWA in general.
@@
Our numerical tests including the O4-CFME method[48, 43] demonstrate superior performance of the ME-based approaches and the benefit of using more Gauss-Legendre points and more exponentials instead of commutators. As yet, these results are limited to the $\Lambda$-model with periodic time dependences as considered here. Thus, further tests are needed for other systems and more complex time dependences in order to establish a reliable general trends of performance. In fact, all of our ME-based propagators are easily applicable to other types of systems[11, 25, 78, 14] that are important in the context of quantum thermodynamics or energy/charge transport. In addition, although we considered only the case of periodic driving, our propagators can easily be applied to non-periodic driving equally well. In this sense, our approaches are complementary to those based on Floquet approaches.[65, 25, 66, 67, 68, 69, 70]
@@
It is also noteworthy to mention the stability of our ME based numerical method for the non-Hermitian dynamics (e.g. Figs. 4). We found that the trace of unity is maintained with only negligible deviations (on the order of $10^{-12}$ in the worst case) due to numerical error, even in the long-time limit. This suggests that, for the time-dependent system Hamiltonian $\hat{H}(t)$ with time-independent Lindblad dissipative terms as in Eq. (15), the ME truncated at any order produces a trace preserving map. In fact, we find numerically that it produces CPTP maps, which not only guarantee that the trace is maintained but ensure that the resulting density operator is physical even if it only describes a subsystem. An analytical proof for the closed system case is shown in Appendix C. For the open system dynamics, the CPTP map is produced by our CFME expression, Eq. (12), at any time step. However, for our expression with commutator, Eq. (11), we confirmed numerically that the CPTP map is preserved only for small enough time step sizes, which ensure that the first order term of the ME is larger than the subsequent terms. This can be understood from the fact that the first order term of the ME defines a CPTP map, while it has been shown that higher order terms contribute to breaking the complete positivity.[44]
@@
SUPPLEMENTARY MATERIAL
The Supplementary Material provides additional calculation results for cases in Table LABEL:table-1 and provides a link to the site where python codes can be accessed.
@@ Acknowledgements.
S.J.J. acknowledges major support of this research from the US Department of Energy, Office of Sciences, Office of Basic Energy Sciences (DE-SC0021413, DE-SC0026114) and partial support during the initial stage of this research from the National Science Foundation (CHE-1900170). S.J.J. also acknowledges travel support from Korea Institute for Advanced Study through its KIAS Scholar program.
@@
Appendix A Derivation of simple commutator free Magnus Expansion propagators
We here provide the derivations of the commutator free propagators defined in Eqs. (3) and (12). For this, let us consider the following generic first order differential equation:
$\frac{dy(t)}{dt}=f(t,y(t))=A(t)y(t),$ (20)
where $y(t)$, $f(t,y(t))$, and $A(t)$ are all operators and we omit explicit operator notation in this section. For the case of dynamics in the Hilbert space, $f(t,y(t))=(-i\hat{H}(t)/\hbar)y(t)$, with $y(t)$ the time evolution operator. For the case of the dynamics in the Liouville space, $f(t,y(t))=-i{\mathcal{L}}(t)y(t)$, with $y(t)$ the density operator. For the above form of the differential equation, application of our 4th order ME propagator[36] results in
$y(t+\delta t)=e^{\delta tc_{0}+\delta t^{3}\frac{1}{12}c_{2}-\delta t^{3}\frac{1}{12}[c_{0},c_{1}]}y(t)+\mathcal{O}(\delta t^{5}),$ (21)
where $c_{n}$ is the $n$th Taylor series coefficient of $A(t)$ and is given by $c_{n}=\frac{1}{n!}\frac{d^{n}}{d\tau^{n}}A(\tau)|_{\tau=t+\delta t/2}$. To develop a commutator free version, we look for a propagation of the following form
$y(t+\delta t)=e^{\delta t(\text{a}_{20}c_{0}+\text{a}_{21}\delta tc_{1}+\delta_{t}^{2}\text{a}_{22}c_{2})}e^{\delta t(\text{a}_{10}c_{0}+\delta t\text{a}_{11}c_{1}+\delta t^{2}\text{a}_{12}c_{2})}y(t)+\mathcal{O}(\delta t^{5}),$ (22)
where $\text{a}_{ij}$ are undetermined coefficients. The stability requirement constrains the coefficients to satisfy $\text{a}_{10}=\text{a}_{20}$, $\text{a}_{11}=-\text{a}_{21}$, $\text{a}_{12}=\text{a}_{22}$. The remaining free coefficients can be determined by applying the BCH formula to Eq. (22), equating the resulting expression with the exponent in Eq. (21), and separating terms by their Taylor series coefficients. This procedure results in the following equations:
$2\delta t\text{a}_{10}c_{0}$ $=$ $\delta tc_{0}$
$2\delta t^{3}\text{a}_{12}c_{2}$ $=$ $\delta t^{3}\frac{1}{12}c_{2}$
$\delta t^{3}\text{a}_{10}\text{a}_{11}[c_{0},c_{1}]$ $=$ $-\delta t^{3}\frac{1}{12}[c_{0},c_{1}]{\color[rgb]{0,0,1}{,}}$ (23)
where equations with $\mathcal{O}(\delta t^{5})$ have been discarded. Solving the above equations gives $\text{a}_{10}=\frac{1}{2}$, $\text{a}_{11}=-\frac{1}{6}$, $\text{a}_{12}=\frac{1}{24}$, which determines all the coefficients. The expression can now be converted to one involving three equally spaced points by using the corresponding Legendre interpolating polynomial for $A(t)$, and substituting in the definition of $c_{n}$. Combining like terms leads to our final commutator free expression:
$y(t+\delta t)=e^{\frac{\delta t}{12}(-A_{k}+4A_{k+1/2}+3A_{k+1})}e^{\frac{\delta t}{12}(3A_{k}+4A_{k+1/2}-A_{k+1})}y(t)+\mathcal{O}(\delta t^{5}),$ (24)
where $A_{k}=A(t_{k})$.
Appendix B Liouville-space transformation
@@
Appendix C Proof that truncated Magnus expansion produces a completely positive trace preserving (CPTP) map
Using the same Liouville space notation as in Appendix B, a map is CPTP if and only if it can be written in the following form:[80]
@@ Appendix C Proof that truncated Magnus expansion produces a completely positive trace preserving (CPTP) map
where $\mathcal{M}_{n}$ is a functional of superoperators in the exponent of Eq. (6), consisting of iterated integrals and commutators (for precise definitions, see Ref. 36) . However, note that because $\hat{H}(t)\otimes\hat{I},\hat{I}\otimes\hat{H}(t)^{T}]=0$, it becomes possible to separate the exponent in Eq. (C) into two terms:
@@
Appendix D Closed system dynamics in the rotating wave frame (RWF)
@@
Appendix E Solution of closed system dynamics using RWA at the TPR conditions
@@
Appendix F Dynamics for periodic Hamiltonian
@@
Appendix G Other 4th order methods
G.1 Runge-Kutta and Adams-Moulton methods
We here describe well-known 4th order algorithms, RK4[75] and AM4,[76] for the case of operator differential equation given by Eq. (20), by adopting those for ordinary differential equations.[75, 76]
@@ G.1 Runge-Kutta and Adams-Moulton methods
G.2 Optimized commutator free Magnus Expansion
@@
In the above equation, $A_{i}$ is shorthand for $A(t+\delta tc_{i})$ where $c_{i}$ is the ith Gauss-Legendre point defined as $c_{1}=\frac{1-\sqrt{3/5}}{2}$ , $c_{2}=\frac{1}{2}$ , $c_{3}=\frac{1+\sqrt{3/5}}{2}$, and the coefficient $a_{ij}$ denotes an element of the following matrix $a$: